# Probing Massive Fields with Multi-Band Gravitational-Wave Observations

### v1

@@ V Conclusion and Discussion
In this work, we investigate the prospect of probing massive fields with space borne detectors that target GWs in the mHz band. We consider the dipole radiation from generic massive spin-0 and spin-1 fields, calculate their imprints on the inspiral waveform and estimate the detachability on the fields. We demonstrate the detectability on the effective charge difference $|\Delta q|$ of LISA by calculating the Fisher information matrix. We find that LISA can constrain $|\Delta q|$ down to $10^{-6}$ for fields with mass below $3\times 10^{-17}{\rm eV}$, given a 4-year observation of a GW150914-like binary.
@@
We further emphasis the implication of multi-band observations on probing new massive fields. We show that multi-band observation of a GW150914-like binary can improve the detectability on $\Delta q$ by at most 3 orders of magnitude for spin-0 and spin-1 fields with mass ranging from $10^{-16}{\rm eV}$ to $10^{-15}{\rm eV}$. Multi-band observation can further distinguish the spin of the fields in such mass range, where the spin of the field cannot be identified with single-band observations even if a non-trivial $\Delta q$ is detected.
@@
Our multi-band detection strategy mainly depends on the measurement of merger time, which could in principle be affected by other beyond GR effects, such as modification of GW propagation speed. We expect such degeneracy can be break by considering multiple multi-band events, because effects from modification of GW speed depends on the distance of the sources while effects from massive fields should depends on the intrinsic parameters of the merger binaries. We shall leave this topic for future investigation.
@@ Acknowledgements.
We thank Lijing Shao for helpful discussions. J. Z. is supported by the scientific research starting grants from University of Chinese Academy of Sciences (grant No. 118900M061), the Fundamental Research Funds for the Central Universities (grant No. E2EG6602X2 and grant No. E2ET0209X2) and the National Natural Science Foundation of China (NSFC) under grant No. 12147103.
@@ Appendix A Corrections on GW phase
For spin-0 field, we have
@@
For spin-1 field, we have

### v3

@@ V Conclusion and Discussion
In this work, we investigate the prospect of probing massive fields with space-borne detectors that target GWs in the millihertz band. We consider the dipole radiation from generic massive spin-0 and spin-1 fields, calculate their imprints on the inspiral waveform and estimate the detectability on the fields. We demonstrate the detectability on the effective charge difference $|\Delta q|$ of LISA by calculating the Fisher information matrix. We find that LISA can constrain $|\Delta q|$ down to $10^{-6}$ for fields with mass below $3\times 10^{-17}\,{\rm eV}$, given a 4-yr observation of a GW150914-like binary. Note that we use GW150914-like event as a benchmark to demonstrate the sensitivity of $|\Delta q|$, while the sensitivity does depend on the binary parameters. For fields with $f_{\mu}>1{\rm Hz}$, the sensitivity is mainly determined by the number of orbital cycles observed in the LVK band, and hence will get improved if the binary has a smaller total mass. Constraints on fields with $10^{-1}{\rm Hz}<f_{\mu}<1{\rm Hz}$ replies on the precision of merger time measured in the millihertz band, which distributes from $0.1$s to $100$s and peaks around $3$[67]. As the precision of merger time mostly depends on the signal-to-noise ratio, the sensitivity of $|\Delta q|$ can be improved if the event has a larger signal-to-noise ratio in the millihertz band. For fields with $f_{\mu}<10^{-1}\,{\rm Hz}$, the sensitivity is mainly determined by the number of orbital cycles observed in the millihertz band. Given a certain observation time, the sensitivity will be improved if the binary has smaller total mass. In practice, the massive field will be probed by stacking multiple events. The sensitivity of $|\Delta q|$ given multiple multiband observations will be investigated in future work. In addition, we assume circular orbits in our analysis, which is usually the case in the LVK band, but is not necessarily true in the millihertz band. Depending on the formation channel, stellar mass binaries observed in the millihertz band could have nonzero eccentricity, for example, see Refs. [68, 69] and the references therein. Since both scalar and vector radiations tend to circularize binaries [68], the radiation power and hence $|\Delta q|$ could be probed by measuring the eccentricity distribution, if the field mass is relatively small such that the radiation turns on in the millihertz band. For fields with larger mass, one cannot probe the fields with eccentricity, which is washed out by GW radiation anyway.
As mentioned in Sec. I, a direct coupling to a spin-0/spin-1 field is constrained by other experiments. For example, precision measurements of binary pulsars impose a constraint of $|\Delta q|<10^{-3}$ for fields with mass lower than $10^{-16}\,{\rm eV}$ [39]. Also, the direct coupling can, in principle, lead to a fifth force, which is constrained by local gravity tests [70]. If we assume the charge difference is of the same order of the charges, the Eöt-Wash experiment indicates a constraint of $|\Delta q|<10^{-1}$ for fields with mass lower than $10^{-3}\,{\rm eV}$ (see, for example, Ref. [71] for the accurate plot of the constraints). One should note that the coupling strength is highly model dependent. It can vary in different systems even assuming the same theory. Therefore, constraints obtained in one experiment does not directly apply to the others.
@@
We further emphasis the implication of multiband observations on probing new massive fields. We show that multiband observation of a GW150914-like binary can improve the detectability on $\Delta q$ by at most 3 orders of magnitude for spin-0 and spin-1 fields with mass ranging from $10^{-16}$ to $10^{-15}$${\rm eV}$. Multiband observation can further distinguish the spin of the fields in such mass range, where the spin of the field cannot be identified with single-band observations even if a nontrivial $\Delta q$ is detected.
@@
Our multiband detection strategy mainly depends on the measurement of merger time, which could, in principle, be affected by other beyond GR effects, such as modification of GW propagation speed. We expect such degeneracy can be broken by considering multiple multiband events, because effects from modification of GW speed depends on the distance of the sources while effects from massive fields should depend on the intrinsic parameters of the merger binaries. We shall leave this topic for future investigation.
@@ Acknowledgements.
We thank Lijing Shao and Yuetong Zhao for helpful discussions. J. Z. is supported by the scientific research starting grants from University of Chinese Academy of Sciences (Grant No. E4EQ6604X2), the Fundamental Research Funds for the Central Universities (Grants No. E2EG6602X2 and No. E2ET0209X2), and the National Natural Science Foundation of China (NSFC) under Grants No. 12147103 and No. 12347103.
@@ Appendix A Corrections on GW phase
where $\Psi(f)=2\pi ft-\phi-\tfrac{\pi}{4}$ is the GW phase and is calculated under the stationary phase approximation,
@@
For the spin-0 field, we have
@@
For the spin-1 field, we have